OpenStudy (fatboi96):

Choose the equation that represents the graph below: Graph of a line passing through points negative 2 comma 0 and 0 comma negative 4 y = −2x − 4 y = 2x + 4 y = −2x + 4 y = 2x − 4

OpenStudy (shansha1):

so it passes through the points of (-2,0) and (0,-4) well if we use slope-intercept form we already have the y intercept of -4 so the default equation is already y=mx-4 so B and C are out to find the slope use the equation of m=(y2-y1)/(x2-x1) which are from the two points m=(-4-0)/(0+2) m=-4/2 m=-2 now we substitute the slope into the equation we partially had to get y=-2x-4

OpenStudy (fatboi96):

thanks

OpenStudy (shansha1):

ur not welcome